$SPAD/src/input matrox.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Screen setup for graphics demos for the matrox card screen width

--)lo clip grdef viewdef draw2d )update
--)lo grimage view2D )up

-- 2D things
   -- demo of adaptive plotting
--)cl all
--)r adaptive
-- these should be set already
setAdaptive true -- this is the default
viewDefaults()   -- ditto
--)r "/spad/graphics/oldInput/curl"
draw(curve(sin(t)*sin(2*t)*sin(3*t),sin(4*t)*sin(5*t)*sin(6*t)),t=0..2*pi())
curl1:=getGraph(c1:=%,1)
setAdaptive false
viewXPos 410
pointColor pastel yellow();
lineColor pastel blue();
--)r "/spad/graphics/oldInput/curl"
rc := range curl1
--draw2D(rc.xRange,rc.yRange,sin(t)*sin(2*t)*sin(3*t),_
--       sin(4*t)*sin(5*t)*sin(6*t),0..2*pi())
draw(curve(sin(t)*sin(2*t)*sin(3*t),sin(4*t)*sin(5*t)*sin(6*t)),_
     t=0..2*pi(),[ranges([rc.0,rc.1])])
curl2:=getGraph(c2:=%,1)
curl:=viewport2D()
title(curl,"Adaptive vs Non-Adaptive")
putGraph(curl,curl1,1)
putGraph(curl,curl2,2)
key curl
curly:=makeViewport2D curl
key curly
resize(curly,368,428)
move(curly,905,588)
-- reset defaults
setAdaptive true
viewDefaults()

   -- implicit graphs
--pointColor pastel red()
--lineColor light blue()
--)r cart
--cart := %
--resize(cart,277,180)
--move(cart,600,483)
viewDefaults()

-- a few 3D things
--)r wiggle
--wiggle:=%
wiggle:=draw(cos(x*y),x=-3..3,y=-3..3)
move(wiggle,876,200)
--)r lumpy
--lumpy:=%
lumpy:=sin(2*x**2 + 3 * y**2)/(x**2 + y**2)
resize(lumpy,300,330)
move(lumpy,600,688)
--)r torus
--donut:=%
fcolor(x:SF):SF == x
--donut:=newdraw(sin t,cos t,0,fcolor,t=0..2*%pi)
donut:=draw(surface(sin t,cos t,0),fcolor,t=0..2*%pi)
resize(donut,374,180)
move(donut,900,0)

———-

References^{[1] nothing}

5